OpenStudy (anonymous):

Divide: (-5/9) ÷(-5/6)= 5/9x6/5=30/45divided by 15=2/3? right...

OpenStudy (whpalmer4):

\[\frac{(\frac{-5}{9})}{(\frac{-5}{6})} = \frac{-5}{9}*\frac{6}{-5}\] because dividing by a fraction is like multiplying by its reciprocal. Cancel the -5's and you're left with 6/9 = 2/3...

OpenStudy (anonymous):

Ty yours is a much simpler way in doing it

OpenStudy (whpalmer4):

@bonjovifan03 if you're confident about not getting mixed up, you could cancel the -5's directly, and just flip the fraction over so the denominator of the denominator becomes the new numerator :-)